SDJ, Inc., has net working capital of $1,910, current liabilities of $5,610, and inventory of $1,265.
Requirement 1:
What is the current ratio? (Round your answer to 2 decimal places (e.g., 3.16).)
  Current ratio times  
Requirement 2:
What is the quick ratio? (Round your answer to 2 decimal places (e.g., 3.16).)
  Quick ratio times  
Answer #1

net working capital=current assets-current liabilities

current assets=(1910+5610)=7520

a.Current ratio=current assets/current liabilities

=7520/5610

=1.34(Approx).

b.Quick ratio=(current assets-inventory)/current liabilities

=(7520-1265)/5610

=1.11(Approx).